The length of a rectangle is 3 m longer than its width. If the perimeter of the rectangle is 38 m find its area
Found 2 solutions by midwood_trail, nerdybill:
Answer by midwood_trail(310)

The length of a rectangle is 3 m longer than its width. If the perimeter of the rectangle is 38 m. Find its area.
Let width = x
Let length = x + 3
The formula to find the perimeter of a rectangle is given to be:
P = 2L + 2W
The perimeter is 38 m.
38 = 2(x + 3) + 2(x)
38 = 2x + 6 + 2x
38 = 4x + 6
38 - 6 = 4x
32 = 4x
32/4 = x
8 = x
I just found x to be 8.
We replace x and x + 3 with 8 and simplify.
Width = x = 8 m
Length = x + 3 = 8 + 3 = 11 m
The width is 8 m and the length is 11 m.
=============================================
However, the question is asking for the area.
Since we now know the width and length, we can use the formula for finding the area of a rectangle, which is A = LW.
Area = 11 m times 8 m
Area = 88m^2
We need to place a (^2) at the end of each area question because area is measured in squared units.
The final answer is:
Area = 88m^2....This is read: "eighty-eight meters squared."
Did you follow?

Answer by nerdybill(7384)

Let x = width
then because "length of a rectangle is 3 m longer than its width"
x+3 = length
.
Perimeter of any rectangle:
2(width + length)
.
so:
38 = 2(x + x+3)
38 = 2(2x+3)
38 = 4x+6
32 = 4x
8 m = x (width)
.
length:
x+3 = 8+3 = 11 m (length)
.
Area of any rectangle:
width * length = 8 * 11 = 88 square meters (answer)